OpenStudy (anonymous):

I need help with this question, im in algebra 1 and stuck can someone please please help. the question is... "Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)"

OpenStudy (anonymous):

(-2x+y=1)..............x2

OpenStudy (anonymous):

-4x+2y=2............(1)

OpenStudy (anonymous):

-4x+y=-1..........(2)

OpenStudy (anonymous):

subtracting eqns 1 and 2,

OpenStudy (anonymous):

I only have 1 and 3 though I have no one and 2